### v1

@@ 1 Introduction
However, the accuracy of current generative AI models’ performance in some tasks such as text summarization or generation, combined with its clarity and confidence of delivery, might create the illusion that it enhances productivity in other domains. Inaccurate, faulty or “hallucinated” output may not be immediately detected and could be used as an input in a knowledge worker’s subsequent production flow. For instance, Kabir et al. (2024) analysed ChatGPT’s response to 517 programming questions and found that 52% of its answers were incorrect and that the users presented with these answers overlooked these errors 39% of the time. Nevertheless, users still tend to prefer to use ChatGPT because of its comprehensive responses (Kabir et al., 2024) and the confident language of the responses (Li et al., 2023).
@@
In this paper, we use observational data to analyse the heterogeneous effects of ChatGPT on the output quantity and quality of experienced and less experienced software developers. In particular, we exploit Italy’s sudden announcement of a ChatGPT ban as a natural experiment to examine the ban’s short-run effects on GitHub users’ productivity. We find that the ban had no systematic effect on the overall output of more experienced developers and only some small negative effects on their output for more routine tasks (resolving issues and debugging). However, among less experienced users, the short-term lack of access to ChatGPT increased both the amount of output and its quality. For this group of users, the likelihood that we observe any output-related activity on GitHub is approximately 10% higher for the two business days following the ban. This effect size shrinks for the subsequent days. In the same vein, we find some tentative evidence that Internet users in Italy adapted fairly quickly to the legislation by increasing their use of virtual private networks (VPNs) and encrypted routing to circumvent the ban. A placebo test shows that our results are unlikely to be driven by unobserved, seasonal factors. The results are also robust to our using alternative outcome variables and analysing the effect at the user–repository–day level.
@@
Our results present some first nonexperimental empirical evidence on the effects of restricting access to generative AI on workers’ performance in more complex tasks. Importantly, we show that the effects of generative AI are heterogeneous by worker skill type.
@@
Our study complements the existing, largely experimental, literature on the effects of generative AI on worker productivity in less complex tasks (e.g., content writing, customer support), where generative AI output is less error-prone, by examining a setting with more complex tasks, where AI-generated output can be less accurate or more faulty. Existing work by Brynjolfsson et al. (2023) and Noy and Zhang (2023) has found mainly positive productivity effects of generative AI and stronger effects for less experienced workers in the contexts of customer support and content writing tasks. In contrast, our results suggest that, for more complex tasks (e.g., code development), [^7] Related work by Peng et al. (2023) and Chatterjee et al. (2024) has found positive effects of GitHub Copilot on both the productivity and job satisfaction of software developers. While GitHub Copilot is also an AI-based tool developed by GitHub and OpenAI, it is specifically a code completion tool. In contrast, ChatGPT is a chat-based (rather than auto-complete) tool and can be used to produce entirely new code segments/programs based on a human language prompt; in such cases, the accuracy of generative AI is highly variable (Kabir et al., 2024). generative AI does not necessarily boost the productivity of less skilled workers and can even decrease their output quantity and quality. Our results, therefore, confirm the findings of Dell’Acqua et al. (2023) in a controlled experiment environment that, for tasks beyond ChatGPT’s current capabilities, using ChatGPT increases the time a worker spends on a task. We complement their results by highlighting the differential effects by knowledge worker skill level. Our finding of heterogeneous effects for more complex tasks also empirically complements the larger discussion in economics on technological change and inequality in the labour market (e.g., Acemoglu, 2002; Autor et al., 2003; Goldfarb and Tucker, 2019; Acemoglu and Restrepo, 2020), in particular the productivity and labour market effects of AI (e.g., Brynjolfsson et al., 2017; Agrawal et al., 2019; Acemoglu, 2021; Eloundou et al., 2023).
@@ 2 ChatGPT and the Italian Ban
ChatGPT is a Generative Pre-trained Transformer (GPT) which is a type of a large language model developed by OpenAI and was publicly launched in November 2022. ChatGPT uses a transformer-based neural network that was trained on the publicly available text corpus from the Internet including digitized books, wikipedia as well as Q&A help forums for software developers such as Stack Overflow. The LLM uses this training data to to identify patterns of words and phrases, predict next words and thereby generate new text as well as code.
@@
On April 1, 2023, the Italian data protection authority (Garante per la protezione dei dati personali) blocked use of the ChatGPT chatbot, citing privacy concerns, and announced an investigation into OpenAI’s compliance with the European Union’s General Data Protection Regulation (GDPR). In particular, the authority stated that there was no legal basis for the mass collection and storage of personal data to train the algorithms underlying the platform’s operation. [^9] Shiona McCallum, “ChatGPT banned in Italy over privacy concerns”, BBC 01/04/2023, https://www.bbc.com/news/technology-65139406 The ban was lifted in late April after OpenAI responded to the data protection authority’s privacy concerns. [^10] Shiona McCallum, “ChatGPT accessible again in Italy”, BBC 28/04/2023, https://www.bbc.com/news/technology-65431914
@@ 3 Data
We access individual-level, real-time activity data for GitHub users in Italy (treatment) and Austria, France, and Spain (control) in the week prior to and that immediately after the ChatGPT ban in Italy (March 27–April 11, 2023).
@@
We access data on individual GitHub users’ activity from the GitHub archive, which is updated daily and contains all public event data. The GitHub archive data are hosted on Google’s BigQuery warehouse system and can be accessed with a query on Google’s cloud infrastructure. GitHub user information such as the year of GitHub user account creation was downloaded with the GitHub GraphQL API. [^13] The Python scripts written to access the GitHub user information for the respective countries are available at https://GitHub.com/sodalabsio/GitHub_scrape. The two datasets are merged via the unique GitHub user login.
We use the individual-level action data to construct three sets of outcome variables: The first group is labelled Output quantity and includes aggregate Output limited to “productive” actions, aggregate Output as defined by Shen (2023), Commits, and Issues closed. The second group, Output quality, contains the Bug fix ratio, Pull requests (PR) merged and the PR merge ratio. The third group, Task complexity, consists of Files edited per PR and Lines added per PR. A detailed description of the construction and definition of each variable is provided in Appendix Table A.1.
On a daily basis, these actions are relatively rare events at the user level. Hence, we transform each into a binary indicator that takes 1 if one of the actions in a category is recorded for the user on a given day and takes 0 otherwise. [^14] The distribution of day–user-level counts of the main event variables during the sample period is presented in Figure B.1 in the appendix. Descriptive statistics at the user–day and user level are presented in Appendix Table B.1.
@@
We compile a list of packages hosted on GitHub for ten analytical programming languages: C, C++, Go, Java, JavaScript, Julia, Perl, Python, R, and Rust. We rely in the first instance on the community–curated “Awesome Lists” to locate GitHub repositories for “popular” packages in each language. In a second step, we scrape the information on all packages hosted on the official software repositories for Python (pypi), R (CRAN) and Julia (JuliaRegistries) to retrieve information on each package’s GitHub repository. We make use of the standardized GitHub URL structure to identify the owner and name of a package repository. [^15] The stylized URL for a package repository is https://GitHub.com/[account name hosting the repository]/[repository name] (for instance, https://github.com/numpy/numpy). To identify the GitHub user accounts other than the owner that contribute to a package repository, we use information on each individual GitHub user’s activity from January 2011 until March 2023. We restrict the list of GitHub event types to “productive” events to select primarily accounts that made at least one substantial contribution to a package repository. [^16] Our set of “productive” event types comprises PullRequestEvent, PullRequestReviewEvent, PullRequestReviewCommentEvent, PushEvent, and ReleaseEvent. Moreover, we winsorise the sample at the 1st and 99th percentiles to safeguard against outliers. [^17] Note that we exclude bot accounts from the list of contributors prior to winsorising. Our final list of package contributors and owners comprises $483,855$ unique GitHub user accounts, of which $5,920$ are part of our baseline sample.
@@ 4 Effect of the ChatGPT Ban on GitHub Output
To analyse the effect of the Italian ChatGPT ban on GitHub users’ output, we estimate variants of the following difference-in-difference (DID) event-study model:
$Y_{it}=$ $\alpha_{i}+\lambda_{t}+\gamma_{dow}+\sum_{\tau=-4}^{-2}\beta_{\tau}D_{it}^{\tau}+\sum_{\tau=0}^{3}\beta_{\tau}D_{it}^{\tau}+\epsilon_{it},$ (1)
where $Y_{it}$ denotes the outcome variable, which, in the baseline analysis, is one of the nine output and task variables. $D$ is a dummy variable equalling one for observations in the treatment group at event–day $\tau$ and zero otherwise, with $\tau=-1$ as the reference period; $\alpha_{i}$ is a vector of user-specific fixed effects; $\lambda_{t}$ is day (date) fixed effects; $\gamma_{dow}$ is day-of-the-week fixed effects; and $\epsilon_{it}$ is the error term. The parameters of interest are $\beta_{\tau}$. We cluster standard errors at the user level.
@@ 4.1 Baseline Results
Table 1 presents the main results. The upper panel shows the estimated coefficients for $D$ for each of the nine output variables, four different samples (overall, less experienced, experienced and package contributors) and a time period covering the last two working days (March 27–28, 2023) before and the first two working days after the ban took effect (April 3–4, 2023). We find post-ban increases for the Italian users on most of our output quantity, output quality and task complexity measures. For example, the aggregate output measure increases by 0.012. At a mean of 0.23, this corresponds to an increase in the average likelihood of an output-relevant GitHub event of approximately 5%. [^18] Refer to Table B.1 for summary statistics and Tables B.2 and B.3 for an extended version of Table 1 containing, among others, the dependent variable mean for each regression. These effects are, however, driven only by the less experienced users, whose aggregate output increases by 0.024, which translates to a 10% increase in the average likelihood of our observing any output action on GitHub. In contrast, the estimated coefficients for experienced users and package contributors are close to zero and not statistically significant, with the notable exception of the Bug fix ratio for package contributors. Here, we find a statistically significant, negative effect of the ban. Taken together, the results suggest that the ban actually increased the productivity of less experienced coders. After the ban took effect, both the amount and quality of their output increased, and they tended to edit more files and write more lines of code. This supports the idea that, prior to the ban, ChatGPT might have had a disruptive effect on the output quantity and quality of less experienced coders. Code produced by ChatGPT might not have compiled, might have required a lot of time to debug or might have distracted less experienced coders in other ways. On the other hand, the ban did not systematically affect experienced coders’ output.
In the lower panel, we reestimate the specifications from the upper panel with a longer pre- and post-treatment period. Overall, the pattern of the results remains qualitatively stable. Quantitatively, the effect size for some of the output measures decreases, indicating that users might have found ways to adapt to the new situation or circumvent the ban. In addition, we now find a systematic decrease in experienced users’ and package contributors’ likelihood of closing issues, consistent with proficient users using ChatGPT mainly for debugging purposes. Figure 1 presents the estimated event-study coefficients $\beta_{\tau}$ from specification 1 for the outcomes of interest for each user type. Importantly, a joint F-test of whether all coefficients on the preevent relative time indicators are jointly zero cannot be rejected at conventional levels for any outcome, alleviating concerns about preexisting trends. Moreover, the event-study results reveal that the treatment effect for less experienced users did indeed peak two days after the ban.

### v2

@@ 1 Introduction
However, the accuracy of current generative AI models’ performance in some tasks such as text summarization or generation, combined with its clarity and confidence of delivery, might create the illusion that it enhances productivity in other domains. Inaccurate, faulty or “hallucinated” output may not be immediately detected and could be used as an input in a knowledge worker’s subsequent production flow. For instance, Kabir et al. (2024) analysed ChatGPT’s response to 517 programming questions and found that 52% of its answers were incorrect and that the users presented with these answers overlooked these errors 39% of the time. Nevertheless, users still tend to prefer to use ChatGPT because of its comprehensive responses (Kabir et al., 2024) and the confident language of the responses (Li et al., 2023). The biggest online discussion and help forum for developers, stack overflow, has banned the use of LLM generated content on the forum because the rate of “getting correct answers” from these tools, it too low. [^5] https://meta.stackoverflow.com/questions/421831/policy-generative-ai-e-g-chatgpt-is-banned
@@
In this paper, we use observational data to analyse the heterogeneous effects of ChatGPT on the output quantity and quality of experienced and less experienced software developers. In particular, we exploit Italy’s sudden announcement of a ChatGPT ban as a natural experiment to examine the ban’s short-run effects on GitHub users’ productivity. We find that the ban had no systematic effect on the overall output of more experienced developers and only some small negative effects on their output for more routine tasks (resolving issues and debugging). However, among less experienced users, the short-term lack of access to ChatGPT increased both the amount of output and its quality. For this group of users, the likelihood that we observe any output-related activity on GitHub is approximately 10% higher for the two business days following the ban. This effect size shrinks for the subsequent days. In the same vein, we find some tentative evidence that Internet users in Italy adapted fairly quickly to the legislation by increasing their use of virtual private networks (VPNs) and encrypted routing to circumvent the ban.
Finding high-frequency and consistently measured, observational data on the output of knowledge workers that is comparable across countries is challenging. We follow a the approach from a number of recent papers (e.g., McDermott and Hansen, 2021; Holub and Thies, 2023; Shen, 2023) that have already exploited the temporal granularity of GitHub activity data and used this data as a proxy for software developers’ productivity. When using this data to proxy productivity, we rely on a number of assumptions which we test in a subsequent robustness section. We show that the results are not driven by changes in working hours or the level of complexity of the task. Considering that the ban was implemented close to a major holiday in the treatment and control countries, we also show that the changes in behaviour are not driven by seasonal factors. Our results are also robust to the use of a set of alternative outcome variables and also analysing the effect at the user–repository–day level.
@@
Our results present some first, nonexperimental empirical evidence on the effects of restricting access to generative AI on workers’ performance in more complex tasks. Importantly, we show that the effects of generative AI are heterogeneous by worker’s experience.
@@
Our study complements the existing, largely experimental, literature on the effects of generative AI on worker productivity in less complex tasks (e.g., content writing, customer support), where generative AI output is less error-prone, by examining a setting with more complex tasks, where AI-generated output can be less accurate or more faulty. Existing work by Brynjolfsson et al. (2023) and Noy and Zhang (2023) has found mainly positive productivity effects of generative AI and stronger effects for less experienced workers in the contexts of customer support and content writing tasks. In contrast, our results suggest that, for more complex tasks (e.g., code development), [^8] Related work by Peng et al. (2023) and Chatterjee et al. (2024) has found positive effects of GitHub Copilot on both the productivity and job satisfaction of software developers. While GitHub Copilot is also an AI-based tool developed by GitHub and OpenAI, it is specifically a code completion tool. In contrast, ChatGPT is a chat-based (rather than auto-complete) tool and can be used to produce entirely new code segments/programs based on a human language prompt; in such cases, the accuracy of generative AI is highly variable (Kabir et al., 2024). generative AI does not necessarily boost the productivity of less experienced workers and can even decrease their output quantity and quality. Our results, therefore, confirm the findings of Dell’Acqua et al. (2023) in a controlled experiment environment that, for tasks beyond ChatGPT’s current capabilities, using ChatGPT increases the time a worker spends on a task. We complement their results by highlighting the differential effects by knowledge worker’s level of experience. Our finding of heterogeneous effects for more complex tasks also empirically complements the larger discussion in economics on technological change and inequality in the labour market (e.g., Acemoglu, 2002; Autor et al., 2003; Goldfarb and Tucker, 2019; Acemoglu and Restrepo, 2020), in particular the productivity and labour market effects of AI (e.g., Brynjolfsson et al., 2017; Agrawal et al., 2019; Acemoglu, 2021; Eloundou et al., 2023).
@@ 2 ChatGPT and the Italian Ban
ChatGPT, an LLM created by US startup OpenAI, has been used by millions of people since it launched in November 2022. Trained on a vast corpus of text data from the Internet as it was in 2021, this large-scale AI language model uses a transformer-based neural network to process natural language. During the training process, the model learned to identify patterns and relationships between words, phrases, and sentences, enabling it to generate text.
@@
On April 1, 2023, the Italian data protection authority (Garante per la protezione dei dati personali) blocked use of the ChatGPT chatbot, citing privacy concerns, and announced an investigation into OpenAI’s compliance with the European Union’s General Data Protection Regulation (GDPR). In particular, the authority stated that there was no legal basis for the mass collection and storage of personal data to train the algorithms underlying the platform’s operation. [^10] Shiona McCallum, “ChatGPT banned in Italy over privacy concerns”, BBC 01/04/2023, https://www.bbc.com/news/technology-65139406 There are a number of reasons to expect that the ban did not have an effect on software developers. During the ban access to other Natural Language Processing (NLP) powered tools that are powered by Open AI, such as GitHub CoPilot, was not affected. In contrast to ChatGPT which was trained on the universe of available online text corpus, GitHub Copilot was specifically trained only on code-repositories and build to produce context-aware codes. It provides more accurate results and is also widely used by code-developers. It is important to note, that our study only estimates the effects of a ban of ChatGPT on developer’s output and not a ban on all LLM based coding support tools. Alternative generative AI tools such as Claude and Bard which were launched a few weeks prior to the ban were still accessible. There were also multiple ways for programmers to immediately circumvent the ban (i.e. VPN) and guides on how to do that where shared online. [^11] https://www.programmareinpython.it/blog/chatgpt-bloccato-in-italia-che-fare/
However, anecdotal evidence shows that shortly after the inception of the ban, Italian developers went online and complained about the disruption caused by restricting access to ChatGPT, “[…] a tool that has become an essential part of [their] daily routine” as software developers. [^12] Semero, Anto “ChatGPT Banned in Italy: Mamma Mia! What’s Going On?” https://medium.com/@antonellosemeraro/chatgpt-banned-in-italy-mamma-mia-whats-going-on-97c44284e331, April 2 2023. Concerns were raised that restricting access to ChatGPT during a time when the field of software development is increasingly fast-paced, poses a severe threat to the competitiveness Italian developers and businesses. [^13] https://www.programmareinpython.it/blog/chatgpt-bloccato-in-italia-che-fare/ Whether the ban was effective or not is ultimately an empirical question, which this paper aims to shed light on.
The ban was lifted in late April after OpenAI responded to the data protection authority’s privacy concerns. [^14] Shiona McCallum, “ChatGPT accessible again in Italy”, BBC 28/04/2023, https://www.bbc.com/news/technology-65431914
@@ 3 Data
We access individual-level, real-time activity data from the GitHub Archive for GitHub users in Italy (treatment) and Austria, France, and Spain (control) in the week prior to and that immediately after the ChatGPT ban in Italy (March 27–April 9, 2023). [^17] We chose these three countries because all of them are part of the European Union and share a common land or sea border. In settings like ours, it is difficult to find objective criteria that help guiding the choice of comparable units. To provide further confidence that our results are not based on the choice of control countries, we conduct a leave-one-out analysis in Figure LABEL:fig:leave-one-out and show that the results are not driven by the composition of the control group. To account for the Easter break starting with Good Friday on 7 April—a public holiday in all four sample countries— we restrict the post-treatment period to Monday 3 April to Thursday 6 April and the corresponding pre-period to Monday 27 March to Thursday 30 March. [^18] For more details on the study’s time line and a graphical illustration please see A in the appendix. GitHub Archive is hosted on Google’s BigQuery warehouse system and contains all public event data of GitHub user, which is updated daily and can be accessed with a query on Google’s cloud infrastructure. GitHub user information such as the year of GitHub user account creation was downloaded with the GitHub GraphQL API. [^19] For more information on how we retrieve GitHub user location information, please see B.1 in the appendix. The two datasets are merged via the unique GitHub user login.
We use the individual-level action data to construct two sets of baseline outcome variables: The first group captures output quantity and quality and includes aggregate Output limited to “productive” actions, aggregate Output as defined by Shen (2023), aggregate Output as defined by Holub and Thies (2023), and the Pull request (PR) merge ratio quantifying how many of a user’s suggested code edits were accepted by the repository (project) owners. The second group gauges task choice and complexity and comprises PR opened, Avg. lines added per opened PR, Avg. lines added per merged PR, Easy issue closed, and Interactive activity. A detailed description of the construction and definition of all outcome variables is provided in appendix Table LABEL:tab:var-defs.
On a daily basis, GitHub actions are relatively rare events at the user level. Hence, we transform each count variable into a binary indicator that equals 1 if one of the actions in a category is recorded for the user on a given day, and 0 otherwise. [^20] The distribution of day–user-level counts of the main event variables during the sample period is presented in Figure B.1 in the appendix. Descriptive statistics at the user–day and user level are presented in appendix Table LABEL:tab:summary-user. To use GitHub actions as a measure for developers’ productivity, we need to ensure that the ban did not affect developers’ working hours. While it is not possible to have data on exact working hours for developers, we are still able to check if time spent on GitHub activity during the day has changed between the pre and post period. Using the timestamp data for each individual GitHub action, we are able to show that the distribution of GitHub activities across hours of the day did not change over the period (see Figure LABEL:fig:distribution-work-activity-Italy).
@@
We compile a list of packages hosted on GitHub for ten analytical programming languages: C, C++, Go, Java, JavaScript, Julia, Perl, Python, R, and Rust. We rely in the first instance on the community–curated “Awesome Lists” to locate GitHub repositories for “popular” packages in each language. In a second step, we scrape the information on all packages hosted on the official software repositories for Python (pypi), R (CRAN) and Julia (JuliaRegistries) to retrieve information on each package’s GitHub repository. We make use of the standardized GitHub URL structure to identify the owner and name of a package repository. [^21] The stylized URL for a package repository is https://GitHub.com/[account name hosting the repository]/[repository name] (for instance, https://github.com/numpy/numpy). To identify the GitHub user accounts other than the owner that contribute to a package repository, we use information on each individual GitHub user’s activity from January 2011 until March 2023. We restrict the list of GitHub event types to “productive” events to select primarily accounts that made at least one substantial contribution to a package repository. [^22] Our set of “productive” event types comprises PullRequestEvent, PullRequestReviewEvent, PullRequestReviewCommentEvent, PushEvent, and ReleaseEvent. Moreover, we winsorise the sample at the 1st and 99th percentiles to safeguard against outliers. [^23] Note that we exclude bot accounts from the list of contributors prior to winsorising. Our final list of package contributors and owners comprises $483,855$ unique GitHub user accounts, of which $5,916$ are part of our baseline sample.
@@ 4 Effect of the ChatGPT Ban on GitHub Output
To analyse the effect of the Italian ChatGPT ban on GitHub users’ output, we estimate variants of the following difference-in-difference (DID) specification:
$Y_{it}=$ $\beta D_{it}+\alpha_{i}+\lambda_{t}+\gamma_{dow}+\sigma_{d}\times t+\epsilon_{it},$ (1)
where $Y_{it}$ denotes the outcome variable, i.e. one of the user-specific output and task variables, and $D_{it}$ is the treatment indicator variable equaling 1 for Italian users in the first four work days post the ChatGPT ban. The parameters of interest is $\beta$. Time-invariant differences between users, including ability and experience, are captured by user fixed effects $\alpha_{i}$, while day (date) fixed effects $\lambda_{t}$ account for daily fluctuations in coding output across users. Additionally, we account for differences in working behaviour across week-days via day-of-the week $\gamma_{dow}$ and include control treatment and control group specific time trends to safeguard against differential time-trends in coding activity between Italian Github users and their European peers. [^24] We estimate alternative specifications of our baseline regression model. Including country-specific linear time-trends or excluding time-trends altogether leaves our baseline estimates quantitatively and qualitatively stable (Table LABEL:tab:did-wo-trend and Table LABEL:tab:did-ctry-trend in the appendix).
To check for potential pre-trends and investigate how the estimated effect evolves over time, rather than averaging over the whole window as in the generalized DID specification, we also estimate the following event-study specification:
$Y_{it}=$ $\sum_{\tau=-4}^{-2}\beta_{\tau}D_{it}^{\tau}+\sum_{\tau=0}^{3}\beta_{\tau}D_{it}^{\tau}+\alpha_{i}+\lambda_{t}+\gamma_{dow}+\sigma_{d}\times t+\epsilon_{it},$ (2)
where $D$ is a dummy variable equalling one for observations in the treatment group at event–day $\tau$ and zero otherwise; with $\tau=-1$ serving as the (excluded) reference period. In both specifications, we cluster standard errors at the user level.
@@ 4.1 Baseline Results
Table 1 presents our baseline DID model estimates for the coefficient of interest $\widehat{\beta}$. The upper panel reports the average treatment effect of the ChatGPT ban on output quantity and quality for four different samples (overall, less experienced, experienced and package contributors), while Panel B presents baseline estimates for task choice and complexity. We find that the ban has overall no significant effect on the output quantity and quality of Italian users or their choice of tasks with the notable exception of a negative and significant impact on their ability to close issues. These aggregate results, however, disguise important heterogeneity across users.
In Panel B, we provide suggestive evidence that the estimated effect on output quantity and quality of less experienced users is not result of a change in tasks. In particular, we do not see that the number of pull requests send for review (column 1) changes nor their complexity (column 2). Further, restricting the focus to lines added on pull requests that were actually merged (column 3), reveals a significant positive effect. While this increase could reflect less efficient code by Italian users post-ban, our estimates provide no support for the notion that the increase in the PR merge ratio is the result of decrease in merged PR task complexity. Similarly, we do not observe a switch to easier issues (column 4) nor to collaborative engagement instead of coding activity (column 5). [^25] In unreported results, we test if the estimated effect on easy issues is the result of restricting the attention to the resolution of those issues. If we consider any form of engagement with easy issues—i.e. commenting, opening or closing—we find that the probability of engaging with easy issues declines for less experienced users (significant at the 10% level), while no change is observed for more sophisticated users. For experienced users, we find null effects of the ban on task allocation and complexity, while the ban appears to have some effects on package contributors. In particular, the decline in the probability of opening a pull request and resolving easy issues might indicate that package contributors tackle less but harder tasks after the ban.
Figure 1 presents the estimated event-study coefficients $\beta_{\tau}$ from specification 2 for less experienced users. Importantly, a joint F-test of whether all coefficients prior to the ban are jointly zero cannot be rejected at conventional levels for any outcome, alleviating concerns about preexisting trends. Further, the event-study results reveal that the treatment effect for less experienced users peaked two days after the ban. Figure LABEL:fig:event-study-exp and LABEL:fig:event-study-pkg in the appendix present the corresponding event-study estimates for experienced users and package contributors, respectively. Results for this subset of users reveal that point estimates are rather volatile over time, particularly for package contributors, and reveal no clear pattern except for issue closed. Overall, our event-study results provide support for our main DID findings that the ban actually increased the productivity of less experienced coders, while more sophisticated users were largely unaffected.